OpenStudy (anonymous):

sec =6. where feta lies in quadrant 4. What is tan?

OpenStudy (anonymous):

Tan is negative. And whats feta? lol

Directrix (directrix):

1 + tan squared theta = sec squared theta tan squared theta = 36 - 1 tan theta = - √ 35

OpenStudy (campbell_st):

4th quadrant cos is positive so 1/cos or sec is positive so cos = 1/6 draw a triangle and find the missing side using pythagoras  6^2 - 1^1 = x^2 \[x = -\sqrt{?}\] so \[\tan (\theta) = 1/\sqrt{35}\]